The dimensions of a door are 3 ft wide by 7 ft 6 in high. If the same amount is added to each dimension of the door, the area is increased by 67.5 ft2. Find the dimensions of the new door.
Answer by Alan3354(69443)

The dimensions of a door are 3 ft wide by 7 ft 6 in high. If the same amount is added to each dimension of the door, the area is increased by 67.5 ft2. Find the dimensions of the new door.
-------------
3*7.5 = 22.5 sq ft
(3+x)*(7.5+x) = 22.5 + 67.5 = 90
x^2 + 10.5x - 67.5 = 0
(x + 15)*(x - 4.5) = 0
x = 4.5 feet